A company assigns overhead cost to completed jobs on the basis of 118% of direct labor cost. The job cost sheet for Job 413 shows that $14,092 in direct materials has been used on the job and that $10,600 in direct labor cost has been incurred. A total of 1,200 units were produced in Job 413.


Required:

a. What is the total manufacturing cost assigned to Job 413?

        

b. What is the unit product cost for Job 413?
Answer #1

Answer :

A) $ 12,508

B) $ 31

a) Total MAnufacturing Cost Assigned To Job 413?

Total overheads = Direct labor *118% = 10,600*118% = $12,508

B) Unit Product Cost for Job 413?

Total Cost / Total Units

Total Cost = 14,092+10,600+12,508 = 37,200

Units Prodcut Cost = $37,200/ 1200 = $31